### v2

# MSCloudCAM: Cross-Attention with Multi-Scale Context for Multispectral Cloud Segmentation

Abstract
Clouds remain a critical challenge in optical satellite imagery, hindering reliable analysis for environmental monitoring, land cover mapping, and climate research. To overcome this, we propose MSCloudCAM, a Cross-Attention with Multi-Scale Context Network tailored for multispectral and multi-sensor cloud segmentation. Our framework exploits the spectral richness of Sentinel-2 (CloudSEN12) and Landsat-8 (L8Biome) data to classify four semantic categories: clear sky, thin cloud, thick cloud, and cloud shadow. MSCloudCAM combines a Swin Transformer backbone for hierarchical feature extraction with multi-scale context modules—ASPP and PSP—for enhanced scale-aware learning. A Cross-Attention block enables effective multisensor and multispectral feature fusion, while the integration of an Efficient Channel Attention Block (ECAB) and a Spatial Attention Module adaptively refine feature representations. Comprehensive experiments on CloudSEN12 and L8Biome demonstrate that MSCloudCAM delivers state-of-the-art segmentation accuracy, surpassing leading baseline architectures while maintaining competitive parameter efficiency and FLOPs. These results underscore the model’s effectiveness and practicality, making it well-suited for large-scale Earth observation tasks and real-world applications.
I Introduction
Clouds significantly obscure the Earth’s surface in optical satellite imagery, creating persistent challenges for remote sensing applications. Accurate detection and classification of cloud types are essential for atmospheric correction, land surface monitoring, and environmental modeling.
Current deep learning–based cloud masking and segmentation methods have demonstrated strong performance by leveraging hierarchical feature representations. However, most existing approaches remain limited in their ability to generalize across different sensors and spectral configurations. This limitation becomes especially critical in multi-class scenarios that require discriminating between thin clouds, thick clouds, and cloud shadows, where subtle spectral and spatial variations are often misclassified.
Figure 1: High-level architecture of MSCloudCAM. The model integrates a Swin Transformer backbone with ASPP and PSP for multi-scale context, and incorporates a Cross Attention module for robust feature fusion across Sentinel-2 and Landsat-8 imagery.
We propose MSCloudCAM, a novel architecture designed for robust cloud segmentation in multispectral, multi-sensor imagery. The model exploits multi-scale context encoding and cross attention to effectively fuse complementary features from different resolutions and spectral domains, enabling accurate pixel-wise classification across heterogeneous datasets.
Our key contributions are as follows:
  • We develop a hybrid backbone that integrates the Swin Transformer [1] with ASPP [2] and PSP [3] modules, enabling comprehensive multi-scale context aggregation.
  • We introduce a Cross Attention module that enhances feature fusion between high-level semantic and mid-level spatial features, improving robustness in complex cloud scenarios.
  • We demonstrate that MSCloudCAM achieves superior performance on both Sentinel-2 (CloudSEN12) [4] and Landsat-8 (L8Biome) [5] imagery, consistently outperforming strong state-of-the-art segmentation networks while maintaining computational efficiency.
II Related Work
II-A Cloud Segmentation in Remote Sensing
Cloud segmentation in optical satellite imagery has long been studied due to its importance in atmospheric correction, land cover mapping, and environmental monitoring. Early rule-based methods, such as Fmask [6], applied band-specific thresholds and decision trees (e.g., cirrus, TIR bands in Landsat) but often misclassified mixed pixels, thin clouds, and bright surfaces like snow or sand [7].
Spectral index–based approaches leveraged indices such as NDSI, NDVI, and cloud probability measures [8], offering more robustness than simple thresholds but still failing under high spectral variability or semi-transparent clouds.
Later, machine learning classifiers (SVMs, Random Forests, gradient boosting) [9] combined spectral and textural features, improving accuracy but remaining limited by handcrafted inputs and shallow context modeling. These shortcomings motivated the transition to deep learning, which directly learns multi-scale, discriminative features from raw multispectral data.
II-B Deep Learning Approaches
Deep learning has become the dominant paradigm for cloud segmentation in remote sensing, offering substantial improvements in accuracy, robustness, and generalization compared to traditional threshold-based methods.
CNN-based architectures remain foundational in this domain due to their strong spatial feature extraction capabilities and hierarchical representation learning. Encoder–decoder designs such as UNet [10] and its lightweight adaptation UNetMobv2 [11] preserve fine-grained localization via skip connections, making them effective for delineating cloud boundaries. DeepLabV3+ [2] enhances context capture through atrous convolutions and the Atrous Spatial Pyramid Pooling (ASPP) module, enabling multi-scale feature aggregation for detecting clouds of varying thickness. Several CNN variants have been developed specifically for cloud detection, including CDNetv1 [12] and CDNetv2 [13] with spectral index integration and multi-branch architectures, KappaMask [14] incorporating Kappa coefficient–based loss for class imbalance handling, DBNet [15] using dual-branch feature extraction for spatial–spectral fusion, SCNN [16] employing spectral–spatial separable convolutions, MCDNet [17] with multi-channel dense connections, and HRCloudNet [18] designed for high-resolution imagery.
Transformer-based architectures leverage self-attention mechanisms to capture long-range dependencies, a property highly beneficial for clouds that span large spatial extents or exhibit non-local spectral correlations. SegFormer [19] combines the lightweight MixVision Transformer (MiT) backbone with an MLP decoder for efficient global context modeling. Mask2Former [20] unifies instance, semantic, and panoptic segmentation through masked attention, offering strong adaptability to heterogeneous cloud patterns. UPerNet-InternImage [21] integrates a hierarchical transformer backbone (InternImage) with pyramid pooling for multi-scale representation, achieving state-of-the-art results on multiple remote sensing segmentation benchmarks.
Hybrid architectures aim to combine the strengths of CNNs and transformers by using CNN layers for low-level texture and spectral edge extraction, followed by transformer modules for global reasoning. Examples include HRNet–Transformer hybrids and cross-attention fusion models [1], which have been applied to atmospheric feature segmentation.
Foundation model–based segmentation has recently emerged as a promising direction, adapting large-scale pretrained vision transformers such as DINOv2 [22] for remote sensing segmentation. These models are trained on massive, diverse datasets and provide rich, generalizable features. While they have demonstrated strong zero-shot and fine-tuned performance in RGB domains, their adaptation to multispectral or multi-sensor data requires spectral-specific preprocessing and fine-tuning strategies.
Despite these advances, most existing models are trained and evaluated on single-sensor datasets, limiting their ability to generalize across multi-sensor imagery. Moreover, few explicitly integrate multi-scale spectral–spatial features with cross-attention mechanisms tailored for cloud segmentation. Our proposed MSCloudCAM addresses both gaps by combining a Swin Transformer[1] backbone with ASPP[2] and PSP[3] modules for multi-scale context aggregation, a Cross Attention block for sensor fusion, and ECAB[23] and spatial attention[24, 25] for adaptive feature refinement.
II-C Gaps in Current Methods
Despite progress, most existing networks are trained on single-sensor datasets, limiting generalization to multi-sensor scenarios. Additionally, few works explicitly integrate multi-scale spectral-spatial features with cross-attention mechanisms for cloud segmentation. Our MSCloudCAM addresses both these gaps.
III Methodology
III-A Overview
MSCloudCAM is a multispectral semantic segmentation network that integrates a Swin Transformer encoder with dual context modules, a cross-attention fusion mechanism, and a lightweight combined attention refinement block. The model is designed to capture long-range spectral–spatial dependencies while retaining fine-grained local structure. A multi-stage decoder with auxiliary supervision produces the final dense predictions. The overall pipeline is illustrated in Fig. 2.
Figure 2: Architecture of MSCloudCAM. A multispectral Swin Transformer encoder generates hierarchical features. ASPP and PSP modules enrich deep and intermediate representations. These are fused by convolutional cross-attention, refined via channel–spatial attention, and decoded by a multi-stage upsampling path with auxiliary supervision.
III-B Backbone Feature Extraction
Let $\mathbf{X}\in\mathbb{R}^{B\times C_{\text{in}}\times H_{0}\times W_{0}}$ denote the input multispectral image with $C_{\text{in}}$ spectral channels (13 for Sentinel-2, 11 for Landsat-8). The Swin Transformer encoder $\mathcal{E}(\cdot)$ produces a hierarchy of features:
$\{f_{1},f_{2},f_{3},f_{4}\}=\mathcal{E}(\mathbf{X}),\quad f_{i}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}},$ (1)
with channel dimensions $[96,192,384,768]$ and progressively reduced spatial sizes $(H_{i},W_{i})$. Shifted-window self-attention within Swin ensures that $f_{4}$ captures global dependencies, while $f_{1}$$f_{3}$ preserve finer detail.
III-C Multi-Scale Context Modules
To enhance representational richness, we employ two complementary context encoders that operate at different feature hierarchies.
ASPP [2]
Applied to the deepest feature map $f_{4}$, Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
$x_{\text{ASPP}}$ $=\phi_{\text{ASPP}}(f_{4})$
$=\bigoplus_{r\in\{1,6,12,18\}}\sigma\!\big(\text{Conv}_{3\times 3}^{(r)}(f_{4})\big)\;\oplus\;\sigma\!\big(\text{GAP}(f_{4})\big),$ (2)
where $\oplus$ denotes channel-wise concatenation, $r$ specifies the dilation rate, GAP is global average pooling, and $\sigma$ represents the ReLU nonlinearity. This module injects large-scale semantic context while retaining spatial discrimination.
PSP [3]
For the intermediate feature map $f_{3}$, Pyramid Scene Parsing aggregates multi-scale contextual cues by adaptive pooling at several grid scales:
$x_{\text{PSP}}=\phi_{\text{PSP}}(f_{3})=\bigoplus_{s\in\{1,2,3,6\}}\sigma\!\left(\text{Up}\big(\text{Pool}_{s}(f_{3})\big)\right),$ (3)
where $\text{Pool}_{s}$ denotes average pooling to an $s\times s$ grid, followed by projection and bilinear upsampling. This captures scene-level information at progressively coarser scales, which is essential for delineating fine structures such as thin clouds.
III-D Cross-Attention Fusion
The ASPP and PSP outputs are concatenated to form a joint query representation:
$x_{\text{cat}}=[\,x_{\text{ASPP}}\,\|\,x_{\text{PSP}}\,]\in\mathbb{R}^{B\times(C_{a}+C_{p})\times H\times W}.$ (4)
A convolutional multi-head cross-attention module then refines this fused tensor by treating $x_{\text{PSP}}$ as contextual guidance:
$Q$ $=W_{Q}*x_{\text{cat}},$ (5)
$K$ $=W_{K}*x_{\text{PSP}},$ (6)
$V$ $=W_{V}*x_{\text{PSP}},$ (7)
$\text{Attn}(Q,K,V)$ $=\operatorname*{softmax}\!\Big(\tfrac{QK^{\top}}{\sqrt{d_{k}}}\Big)V,$ (8)
$\tilde{x}$ $=W_{O}*\text{Attn}(Q,K,V)+x_{\text{cat}}.$ (9)
Here, * denotes convolution, $d_{k}$ is the per-head dimension, and residual addition ensures stable learning. This mechanism aligns global semantics from ASPP with fine-grained spatial cues from PSP.
III-E Bottleneck Projection
The fused representation $\tilde{x}$ is compressed and regularized through a three-layer bottleneck:
[…]
$))),\displaystyle=\sigma\!\Big(\text{Conv}_{1\times 1}\!\big(\sigma(\text{Conv}_{3\times 3}\!(\sigma(\text{Conv}_{1\times 1}(\tilde{x}$$)\big)\Big)$, (10)
which first reduces dimensionality, then captures local spatial dependencies, and finally projects features into a compact 512-channel embedding. This step reduces redundancy and prepares the representation for attention-based refinement.
III-F Combined Attention Refinement
The bottleneck output is recalibrated by a combined attention operator:
$z^{\prime}=\phi_{\text{CA}}(z)=\phi_{\text{ECA}}(z)\odot\phi_{\text{SA}}(z)+z,$ (11)
where $\phi_{\text{ECA}}$ denotes Efficient Channel Attention [23] and $\phi_{\text{SA}}$ is spatial attention [24, 25], and $\odot$ is element-wise multiplication. This highlights discriminative regions while preserving residual information.
III-G Decoder with Deep Supervision
The refined representation $z^{\prime}$ is progressively upsampled through transposed convolutions:
$y_{i}=\sigma(\text{Deconv}_{i}(y_{i-1})),\quad i=1,\dots,4,$ (12)
with $y_{0}=z^{\prime}$. Two intermediate outputs, $\hat{Y}_{1}$ and $\hat{Y}_{2}$, are generated from $y_{1}$ and $y_{2}$, respectively, and resized to the input resolution $(H_{0},W_{0})$. The final prediction $\hat{Y}$ is obtained from $y_{4}$. The training objective jointly supervises all three outputs:
$\mathcal{L}=\lambda_{\text{final}}\mathcal{L}(\hat{Y},Y)+\lambda_{1}\mathcal{L}(\hat{Y}_{1},Y)+\lambda_{2}\mathcal{L}(\hat{Y}_{2},Y),$ (13)
where $Y$ is the ground truth mask, and $(\lambda_{\text{final}},\lambda_{1},\lambda_{2})$ are supervision weights.
IV Experiments
TABLE I: Performance comparison on CloudSEN12 L1C (TOA).
Model IoU (%) F1 (%) Accuracy (%) aAcc
Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
DBNet [15] 85.20 43.02 81.40 52.44 65.52 92.01 60.16 89.75 68.80 77.68 95.56 53.15 89.40 62.47 75.15 86.83
CDNetv2 [13] 84.98 43.37 80.67 53.40 65.60 91.88 60.50 89.30 69.62 77.83 94.80 53.13 89.32 65.34 75.65 86.68
HRCloudNet [18] 86.01 45.88 83.54 57.61 68.26 92.48 62.90 91.03 73.10 79.88 94.49 59.47 91.39 67.19 78.13 87.86
U-Net-MobV2 90.82 85.98 50.54 67.21 73.64 95.19 92.46 67.14 80.39 83.80 94.42 96.11 95.05 96.62 95.55 91.10
U-Net++-MobV2 90.93 86.58 49.78 66.93 73.56 95.25 92.81 66.47 80.19 83.68 94.47 96.25 95.27 96.56 95.64 91.28
DeepLabV3+-MobV2 90.13 85.75 48.97 66.29 72.79 94.81 92.33 65.75 79.73 83.15 94.02 96.04 94.53 96.51 95.28 90.55
U-Net-ResNet101 90.62 86.16 47.15 66.05 72.50 95.08 92.56 64.08 79.56 82.82 94.24 96.13 94.97 96.57 95.48 90.95
U-Net++-ResNet101 90.90 86.43 46.24 68.46 73.01 95.23 92.72 63.24 81.27 83.12 94.45 96.16 95.06 96.74 95.60 91.20
DeepLabV3+-ResNet101 90.65 85.66 50.60 66.36 73.32 95.10 92.28 67.20 79.78 83.59 94.38 95.94 94.94 96.47 95.43 90.87
UPerNet-InternImage 90.85 85.28 36.70 68.87 70.42 95.09 91.82 47.94 80.98 78.96 94.50 96.43 95.24 96.85 95.75 91.51
Mask2Former-DINOv2 88.50 82.10 43.61 63.62 69.46 93.90 90.17 60.73 77.76 80.64 92.94 94.84 94.41 96.11 94.57 89.15
Mask2Former-Swin-T 90.75 86.70 49.90 69.53 74.22 95.15 92.88 66.58 82.02 84.16 94.46 96.32 94.34 96.93 95.51 91.03
SegFormer-MiT-B5 91.47 86.70 49.77 70.24 74.54 95.55 92.88 66.46 82.52 84.35 94.84 96.27 95.25 96.92 95.82 91.64
MSCloudCAM 91.64 87.24 52.58 70.63 75.52 95.64 93.19 68.92 82.78 85.13 94.93 96.42 95.58 97.04 95.99 91.99
IV-A Datasets
We evaluate MSCloud on two publicly available multispectral cloud segmentation datasets.
CloudSEN12 [4] is a global Sentinel-2 cloud segmentation dataset with 13 spectral bands (10–60 m) and pixel-level labels for clear (0), thick cloud (1), thin cloud (2), and cloud shadow (3). We use the high subset, selecting 10k samples each from L1C and L2A. All bands are normalized by dividing top-of-atmosphere reflectance by $3000.0$.
L8Biome [5] is derived from Landsat-8 CCA, containing 11 bands (30 m) with manual masks. The original five classes—fill (0), shadow (64), clear (128), thin cloud (192), and cloud (255)—are remapped to CloudSEN12’s taxonomy: clear $\rightarrow 0$, thick cloud $\rightarrow 1$, thin cloud $\rightarrow 2$, shadow $\rightarrow 3$. The fill class is ignored. Bands are normalized with the same $3000.0$ divisor for consistency.
TABLE II: Dataset split statistics for CloudSEN12 and L8Biome.
Dataset Train Validation Test
CloudSEN12 L1C 8,500 500 1,000
CloudSEN12 L2A 8,500 500 1,000
L8Biome 13,378 4,459 4,459
IV-B Loss Function
The training objective is a weighted sum of the losses from the final prediction and the two auxiliary predictions:
$\mathcal{L}=\lambda_{\text{final}}\mathcal{L}_{\text{final}}+\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{2}.$ (14)
We set $\lambda_{\text{final}}=1.0$ and $\lambda_{1}=\lambda_{2}=0.4$, ensuring the final prediction dominates while auxiliary supervision stabilizes early-stage training.
IV-C Training Protocol
Our model is trained using the Adam optimizer with an initial learning rate of $1\times 10^{-4}$. The batch size is set to 8, and we train for 100 epochs.
IV-D Evaluation Metrics
Per-class Metrics.
$\mathrm{IoU}_{i}$ $=\frac{\mathrm{TP}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}},$ (15)
$\mathrm{F1}_{i}$ $=\frac{2\,\mathrm{TP}_{i}}{2\,\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}},$ (16)
$\mathrm{Acc}_{i}$ $=\frac{\mathrm{TP}_{i}+\mathrm{TN}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}+\mathrm{TN}_{i}}.$ (17)
Per-class metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes, giving equal weight regardless of class size. In contrast, overall accuracy (aAcc) is computed over all pixels, thus weighting larger classes more heavily.
Ignored labels.
Pixels with sentinel values (e.g., Fill in L8Biome) are removed from $\mathbf{C}$ and all denominators via an ignore_index mask, ensuring metrics are based on identical valid-pixel sets.
Reporting.
We report class-wise IoU and F1 (clear, thick cloud, thin cloud, cloud shadow), together with mIoU, mDice, mAcc, and aAcc. Unless noted, averages are macro over classes.
V Results
V-A Quantitative Performance
We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables IIV present the detailed results.
TABLE III: Performance comparison on CloudSEN12 L2A (BOA).
Model IoU (%) F1 (%) Accuracy (%) aAcc
Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
DBNet [15] 85.42 42.76 80.80 53.62 65.65 92.14 59.90 89.38 69.81 77.81 94.83 53.53 90.27 63.45 75.52 86.82
CDNetv2 [13] 85.35 44.01 80.85 54.00 66.05 92.09 61.12 89.41 70.13 78.19 95.65 56.37 88.06 63.56 75.91 86.88
HRCloudNet [18] 87.24 44.18 82.78 59.20 68.35 93.19 61.28 90.58 74.37 79.85 95.41 51.00 93.17 69.83 77.35 88.35
U-Net-MobV2 88.74 83.87 54.13 60.15 71.72 94.03 91.23 70.24 75.12 82.65 93.45 95.55 93.71 95.43 94.53 89.07
U-Net++-MobV2 87.45 83.33 50.70 57.62 69.78 93.30 90.91 67.29 73.11 81.15 92.52 95.47 93.26 95.23 94.12 88.23
DeepLabV3+-MobV2 88.01 83.70 50.36 61.78 70.96 93.62 91.12 66.98 76.37 82.03 92.93 95.46 93.83 95.45 94.42 88.84
U-Net-ResNet101 86.51 81.76 37.12 58.89 66.07 92.77 89.97 54.14 74.12 77.75 91.83 94.73 92.76 95.13 93.61 87.23
U-Net++-ResNet101 87.85 81.03 47.47 59.12 68.87 93.53 89.52 64.38 74.31 80.44 92.84 94.47 93.39 95.37 94.02 88.04
DeepLabV3+-ResNet101 87.96 82.83 50.36 59.52 70.17 93.59 90.61 66.99 74.63 81.45 92.93 95.15 93.79 95.14 94.25 88.50
UPerNet-InternImage 87.79 82.22 40.31 63.96 68.57 93.30 89.78 51.63 77.19 77.98 93.08 95.55 93.48 95.50 94.40 88.81
Mask2Former-DINOv2 83.53 79.32 45.51 56.50 66.21 91.03 88.47 62.55 72.21 78.56 90.30 94.23 91.17 94.93 92.66 85.31
Mask2Former-Swin-T 88.91 84.35 50.72 63.25 71.81 94.13 91.51 67.30 77.49 82.61 93.61 95.77 93.33 95.30 94.50 89.01
SegFormer-MiT-B5 89.28 83.50 51.87 64.19 72.21 94.34 91.01 68.31 78.19 82.96 93.76 95.36 93.86 95.90 94.72 89.44
MSCloudCAM 89.28 84.30 55.53 63.38 73.12 94.33 91.48 71.40 77.58 83.70 93.76 95.69 94.20 95.74 94.85 89.69
TABLE IV: Performance comparison on L8Biome.
Model IoU (%) F1 (%) Accuracy (%) aAcc
Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
CDNetv2 [13] 73.81 24.85 75.85 00.00 43.63 84.93 39.81 86.27 00.00 52.75 90.18 31.95 89.81 00.00 52.98 78.56
HRCloudNet [18] 72.41 30.89 70.76 00.00 43.51 83.99 47.20 82.88 00.00 53.52 85.50 43.70 85.88 00.00 53.77 77.04
DBNet [15] 80.19 38.18 82.37 04.90 51.41 89.00 55.25 90.34 09.34 60.99 96.51 49.34 87.89 05.03 59.70 83.62
U-Net-MobV2 88.98 78.59 39.97 6.37 53.48 94.17 88.01 57.11 11.98 62.82 94.26 91.50 87.55 98.55 92.97 85.93
U-Net++-MobV2 88.96 81.68 39.44 4.42 53.62 94.16 89.91 56.57 8.47 62.28 94.13 93.55 88.76 98.57 93.55 87.09
DeepLabV3+-MobV2 87.25 79.21 33.37 2.54 50.59 93.19 88.40 50.05 4.96 59.15 93.32 91.25 87.50 98.54 92.65 85.31
U-Net-ResNet101 84.29 77.00 30.56 1.25 48.28 91.47 87.01 46.81 2.48 56.94 91.30 90.94 86.06 98.01 91.58 83.16
U-Net++-ResNet101 88.16 80.23 34.94 3.81 51.78 93.71 89.03 51.78 7.34 60.46 93.73 91.94 87.87 98.52 93.01 86.03
DeepLabV3+-ResNet101 87.56 74.19 34.24 4.88 50.22 93.37 85.18 51.01 9.31 59.72 93.47 89.56 85.51 98.56 91.78 83.55
UPerNet-InternImage 54.68 46.97 18.65 1.56 30.46 60.19 53.40 26.72 2.55 35.72 91.00 90.58 83.73 98.35 90.92 81.83
Mask2Former-DINOv2 69.67 70.75 34.67 13.13 47.06 82.12 82.87 51.49 23.22 59.93 84.03 86.86 84.39 98.65 88.48 76.96
Mask2Former-Swin-T 84.36 78.64 37.50 13.54 53.51 91.52 88.04 54.55 23.85 64.49 91.74 91.62 86.01 98.62 92.00 84.00
SegFormer-MiT-B5 88.62 74.38 39.52 15.27 54.45 93.97 85.31 56.65 26.50 65.61 94.04 90.12 86.05 98.70 92.23 84.45
MSCloudCAM 88.27 82.23 40.19 21.48 58.04 93.77 90.25 57.34 35.36 69.18 93.84 92.91 88.74 98.70 93.55 87.09
V-B Qualitative Results
To further illustrate the effectiveness of the proposed method, we present visual comparisons of cloud segmentation results in Fig. 3. Each row corresponds to a different dataset, and each column corresponds to a different segmentation method. As shown in the figure, the proposed method produces more accurate delineation of thin clouds and cloud shadows, while reducing false detections compared to the baseline approaches.
Figure 3: Qualitative cloud segmentation results. Each row corresponds to a dataset. The proposed method produces sharper delineation of thin clouds and shadows compared to other approaches. Class–color mapping: Class 1 (Clear sky – light blue), Class 2 (Thick cloud – white), Class 3 (Thin cloud – light gray), Class 4 (Cloud shadow – dark gray).
V-C Model Complexity Analysis
Table V summarizes the computational complexity of all evaluated models in terms of FLOPs and parameter counts. While some architectures (e.g., UPerNet-InternImage and Mask2Former-DINOv2) exhibit significantly higher complexity, our proposed MSCloudCAM achieves a favorable trade-off between accuracy and efficiency. Specifically, MSCloudCAM attains competitive performance while requiring fewer FLOPs than SegFormer-MiT-B5, and substantially fewer FLOPs than UPerNet-InternImage, with a moderate parameter count.
TABLE V: Model complexity comparison in terms of FLOPs and parameters.
Model FLOPs (G) Parameters (M)
U-Net-MobileNetV2 13.613 6.629
U-Net++-MobileNetV2 18.023 6.825
DeepLabV3+-MobileNetV2 6.057 4.379
U-Net-ResNet101 62.258 51.514
U-Net++-ResNet101 250.000 67.978
DeepLabV3+-ResNet101 56.192 45.670
UPerNet-InternImage-B-DCNv4 200–250 100–130
Mask2Former-Swin-T 73.090 47.400
Mask2Former-DINOv2-ViT-B/14 75–95 87–92
SegFormer-MiT-B5 55.430 45.652
MSCloudCAM 38.98 47.44
VI Code and Model Availability
The full implementation of our proposed MS-CloudCAM framework is publicly available. The source code, training scripts, and experimental configurations are hosted on GitHub: https://github.com/mazid-rafee/ms-cloudcam. Additionally, pretrained model checkpoints are released on Hugging Face: https://huggingface.co/mazid-rafee/MS-CloudCAM/tree/main.
VII Conclusion
We presented MSCloudCAM, a cross attention with multi-scale context network for multispectral, multi-sensor cloud segmentation. By integrating transformer-based backbones with multi-scale context and attention mechanisms, our method achieves superior segmentation accuracy over strong baselines on CloudSEN12 and L8Biome datasets. Future work will explore lightweight variants for onboard satellite processing and extend the model to spatiotemporal cloud tracking.
Acknowledgment
This material is based in part upon work supported by the National Science Foundation under Grants CNS-2018611 and CNS-1920182.

### v3

@@
-MSCloudCAM: Cross-Attention with Multi-Scale Context for Multispectral Cloud Segmentation
+MSCloudCAM: Multi-Scale Context Adaptation with Convolutional Cross-Attention for Multispectral Cloud Segmentation
 Abstract
-Clouds remain a critical challenge in optical satellite imagery, hindering reliable analysis for environmental monitoring, land cover mapping, and climate research.
-To overcome this, we propose MSCloudCAM, a Cross-Attention with Multi-Scale Context Network tailored for multispectral and multi-sensor cloud segmentation.
-Our framework exploits the spectral richness of Sentinel-2 (CloudSEN12) and Landsat-8 (L8Biome) data to classify four semantic categories: clear sky, thin cloud, thick cloud, and cloud shadow.
-MSCloudCAM combines a Swin Transformer backbone for hierarchical feature extraction with multi-scale context modules—ASPP and PSP—for enhanced scale-aware learning.
-A Cross-Attention block enables effective multisensor and multispectral feature fusion, while the integration of an Efficient Channel Attention Block (ECAB) and a Spatial Attention Module adaptively refine feature representations.
-Comprehensive experiments on CloudSEN12 and L8Biome demonstrate that MSCloudCAM delivers state-of-the-art segmentation accuracy, surpassing leading baseline architectures while maintaining competitive parameter efficiency and FLOPs.
-These results underscore the model’s effectiveness and practicality, making it well-suited for large-scale Earth observation tasks and real-world applications.
+Clouds remain a major obstacle in optical satellite imaging, limiting accurate environmental and climate analysis.
+To address the strong spectral variability and the large scale differences among cloud types, we propose MSCloudCAM, a novel multi-scale context adapter network with convolution based cross-attention tailored for multispectral and multi-sensor cloud segmentation.
+A key contribution of MSCloudCAM is the explicit modeling of multiple complementary multi-scale context extractors.
+And also, rather than simply stacking or concatenating their outputs, our formulation uses one extractor’s fine-resolution features and the other extractor’s global contextual representations enabling dynamic, scale-aware feature selection.
+Building on this idea, we design a new convolution-based cross attention adapter that effectively fuses localized, detailed information with broader multi-scale context.
+Integrated with a hierarchical vision backbone and refined through channel and spatial attention mechanisms, MSCloudCAM achieves strong spectral–spatial discrimination.
+Experiments on various multisensor datatsets e.g.
+CloudSEN12 (Sentinel-2) and L8Biome (Landsat-8) show that MSCloudCAM outperforms recent state-of-the-art models while maintaining competitive model complexity, highlighting the novelty and effectiveness of the proposed design for large-scale Earth observation.
 I Introduction
@@
 Accurate detection and classification of cloud types are essential for atmospheric correction, land surface monitoring, and environmental modeling.
-Current deep learning–based cloud masking and segmentation methods have demonstrated strong performance by leveraging hierarchical feature representations.
-However, most existing approaches remain limited in their ability to generalize across different sensors and spectral configurations.
-This limitation becomes especially critical in multi-class scenarios that require discriminating between thin clouds, thick clouds, and cloud shadows, where subtle spectral and spatial variations are often misclassified.
-Figure 1: High-level architecture of MSCloudCAM.
-The model integrates a Swin Transformer backbone with ASPP and PSP for multi-scale context, and incorporates a Cross Attention module for robust feature fusion across Sentinel-2 and Landsat-8 imagery.
-We propose MSCloudCAM, a novel architecture designed for robust cloud segmentation in multispectral, multi-sensor imagery.
-The model exploits multi-scale context encoding and cross attention to effectively fuse complementary features from different resolutions and spectral domains, enabling accurate pixel-wise classification across heterogeneous datasets.
+Recent deep learning–based cloud masking and segmentation methods have shown strong performance by leveraging hierarchical feature representations.
+However, most existing approaches struggle to generalize across heterogeneous sensors and spectral configurations.
+This limitation becomes especially critical in multi-class cloud segmentation requires modeling subtle spectral differences and large variations in spatial scale.
+Figure 1: High-level architecture of MSCloudCAM. The model integrates a hierarchical vision backbone with dual multi-scale context extractors and introduces a novel cross-attention module.
+To address these challenges, we propose MSCloudCAM, a cross-attention with multi-scale context network designed for robust cloud segmentation in multispectral, multi-sensor imagery.
+MSCloudCAM leverages complementary multi-scale context encoders and a newly designed cross-attention mechanism to dynamically fuse fine-grained spatial features with global contextual information.
+This design enables accurate pixel-wise classification across datasets with differing spectral responses, resolutions, and cloud characteristics.
 Our key contributions are as follows:
-  • We develop a hybrid backbone that integrates the Swin Transformer [1] with ASPP [2] and PSP [3] modules, enabling comprehensive multi-scale context aggregation.
-  • We introduce a Cross Attention module that enhances feature fusion between high-level semantic and mid-level spatial features, improving robustness in complex cloud scenarios.
-  • We demonstrate that MSCloudCAM achieves superior performance on both Sentinel-2 (CloudSEN12) [4] and Landsat-8 (L8Biome) [5] imagery, consistently outperforming strong state-of-the-art segmentation networks while maintaining computational efficiency.
+  • We introduce a hybrid multi-scale context framework that leverages two complementary multi-scale context extractors.
+One emphasizes on fine-grained spatial details and the other on capturing broad global semantics.
+Thus it enables us to achieve comprehensive multi-scale feature utilization capability from a hierarchical vision backbone.
+  • We design a novel convolution-based cross-attention module, i.e.
+CAM that leverages these complementary feature hierarchies through an adaptive formulation.
+This mechanism enables dynamic, scale-aware feature fusion, offering a more principled alternative to conventional stacking or concatenation.
+  • We demonstrate that MSCloudCAM achieves superior performance on both Sentinel-2 (CloudSEN12) [1] and Landsat-8 (L8Biome) [2], consistently outperforming state-of-the-art models while maintaining competitive computational efficiency.
 II Related Work
 II-A Cloud Segmentation in Remote Sensing
-Cloud segmentation in optical satellite imagery has long been studied due to its importance in atmospheric correction, land cover mapping, and environmental monitoring.
-Early rule-based methods, such as Fmask [6], applied band-specific thresholds and decision trees (e.g., cirrus, TIR bands in Landsat) but often misclassified mixed pixels, thin clouds, and bright surfaces like snow or sand [7].
-Spectral index–based approaches leveraged indices such as NDSI, NDVI, and cloud probability measures [8], offering more robustness than simple thresholds but still failing under high spectral variability or semi-transparent clouds.
-Later, machine learning classifiers (SVMs, Random Forests, gradient boosting) [9] combined spectral and textural features, improving accuracy but remaining limited by handcrafted inputs and shallow context modeling.
-These shortcomings motivated the transition to deep learning, which directly learns multi-scale, discriminative features from raw multispectral data.
+Cloud segmentation is essential for atmospheric correction, land cover mapping, and environmental monitoring.
+Early rule-based methods such as Fmask [3] relied on band thresholds and decision trees but frequently misclassified thin clouds, mixed pixels, and bright surfaces.
+Spectral index–based approaches (e.g., NDSI, NDVI, cloud probability) [4] provided improved robustness yet still struggled under strong spectral variability or semi-transparent clouds.
+Traditional machine learning classifiers (SVM, Random Forests, boosting) [5] incorporated spectral–textural cues but remained constrained by handcrafted features and limited context modeling, motivating the shift toward deep learning.
 II-B Deep Learning Approaches
-Deep learning has become the dominant paradigm for cloud segmentation in remote sensing, offering substantial improvements in accuracy, robustness, and generalization compared to traditional threshold-based methods.
-CNN-based architectures remain foundational in this domain due to their strong spatial feature extraction capabilities and hierarchical representation learning.
-Encoder–decoder designs such as UNet [10] and its lightweight adaptation UNetMobv2 [11] preserve fine-grained localization via skip connections, making them effective for delineating cloud boundaries.
-DeepLabV3+ [2] enhances context capture through atrous convolutions and the Atrous Spatial Pyramid Pooling (ASPP) module, enabling multi-scale feature aggregation for detecting clouds of varying thickness.
-Several CNN variants have been developed specifically for cloud detection, including CDNetv1 [12] and CDNetv2 [13] with spectral index integration and multi-branch architectures, KappaMask [14] incorporating Kappa coefficient–based loss for class imbalance handling, DBNet [15] using dual-branch feature extraction for spatial–spectral fusion, SCNN [16] employing spectral–spatial separable convolutions, MCDNet [17] with multi-channel dense connections, and HRCloudNet [18] designed for high-resolution imagery.
-Transformer-based architectures leverage self-attention mechanisms to capture long-range dependencies, a property highly beneficial for clouds that span large spatial extents or exhibit non-local spectral correlations.
-SegFormer [19] combines the lightweight MixVision Transformer (MiT) backbone with an MLP decoder for efficient global context modeling.
-Mask2Former [20] unifies instance, semantic, and panoptic segmentation through masked attention, offering strong adaptability to heterogeneous cloud patterns.
-UPerNet-InternImage [21] integrates a hierarchical transformer backbone (InternImage) with pyramid pooling for multi-scale representation, achieving state-of-the-art results on multiple remote sensing segmentation benchmarks.
-Hybrid architectures aim to combine the strengths of CNNs and transformers by using CNN layers for low-level texture and spectral edge extraction, followed by transformer modules for global reasoning.
-Examples include HRNet–Transformer hybrids and cross-attention fusion models [1], which have been applied to atmospheric feature segmentation.
-Foundation model–based segmentation has recently emerged as a promising direction, adapting large-scale pretrained vision transformers such as DINOv2 [22] for remote sensing segmentation.
-These models are trained on massive, diverse datasets and provide rich, generalizable features.
-While they have demonstrated strong zero-shot and fine-tuned performance in RGB domains, their adaptation to multispectral or multi-sensor data requires spectral-specific preprocessing and fine-tuning strategies.
-Despite these advances, most existing models are trained and evaluated on single-sensor datasets, limiting their ability to generalize across multi-sensor imagery.
-Moreover, few explicitly integrate multi-scale spectral–spatial features with cross-attention mechanisms tailored for cloud segmentation.
-Our proposed MSCloudCAM addresses both gaps by combining a Swin Transformer[1] backbone with ASPP[2] and PSP[3] modules for multi-scale context aggregation, a Cross Attention block for sensor fusion, and ECAB[23] and spatial attention[24, 25] for adaptive feature refinement.
+Deep learning has become the dominant strategy for cloud segmentation due to its ability to model spatial–spectral structure directly from multispectral inputs.
+CNN-based architectures such as UNet [6] and UNetMobv2 [7] capture fine localization through skip connections.DeepLabV3+ [8] enhances multi-scale context via dilated convolutions and ASPP.
+Numerous CNN variants target cloud detection specifically, including CDNetv1/v2 [9, 10], KappaMask [11], DBNet [12], SCNN [13], MCDNet [14], and HRCloudNet [15], each improving spectral–spatial fusion or multi-scale representation.
+Transformer-based methods leverage self-attention for long-range dependency modeling, crucial for large or spatially diffuse clouds.
+SegFormer [16] uses a lightweight MiT backbone for global context, while Mask2Former [17] delivers strong generalization through masked attention.
+UPerNet-InternImage [18] combines hierarchical transformers with pyramid pooling to achieve state-of-the-art results in remote sensing segmentation.
+Hybrid models merge CNNs for low-level edge and texture extraction with transformer blocks for global reasoning, including HRNet–Transformer variants and cross-attention designs [19].
+Foundation-model–based segmentation adapts pretrained vision transformers such as DINOv2 [20] to remote sensing; however, effective use of multispectral data still requires spectral alignment and domain-specific fine-tuning.
 II-C Gaps in Current Methods
-Despite progress, most existing networks are trained on single-sensor datasets, limiting generalization to multi-sensor scenarios.
-Additionally, few works explicitly integrate multi-scale spectral-spatial features with cross-attention mechanisms for cloud segmentation.
-Our MSCloudCAM addresses both these gaps.
+Despite significant progress, most approaches are designed for single-sensor datasets, limiting cross-sensor generalization.
+Moreover, few models explicitly integrate complementary multi-scale spectral–spatial cues using directional cross-attention.
+Our MSCloudCAM addresses both challenges.
 III Methodology
 III-A Overview
-MSCloudCAM is a multispectral semantic segmentation network that integrates a Swin Transformer encoder with dual context modules, a cross-attention fusion mechanism, and a lightweight combined attention refinement block.
+MSCloudCAM is a multispectral semantic segmentation network that integrates a hierarchical vision backbone e.g.
+Swin Transformer [19] encoder with dual context extractor modules, a cross-attention fusion mechanism, and a lightweight combined attention refinement block.
 The model is designed to capture long-range spectral–spatial dependencies while retaining fine-grained local structure.
@@
 Let $\mathbf{X}\in\mathbb{R}^{B\times C_{\text{in}}\times H_{0}\times W_{0}}$ denote the input multispectral image with $C_{\text{in}}$ spectral channels (13 for Sentinel-2, 11 for Landsat-8).
-The Swin Transformer encoder $\mathcal{E}(\cdot)$ produces a hierarchy of features:
+The encoder $\mathcal{E}(\cdot)$ produces a hierarchy of features:
 $\{f_{1},f_{2},f_{3},f_{4}\}=\mathcal{E}(\mathbf{X}),\quad f_{i}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}},$ (1)
@@
 III-C Multi-Scale Context Modules
-To enhance representational richness, we employ two complementary context encoders that operate at different feature hierarchies.
-ASPP [2]
-Applied to the deepest feature map $f_{4}$, Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
+To enhance representational richness, we employ two complementary multi context encoders that operate at different feature hierarchies.
+As a part of dual multiscale context extractor, ASPP [8] is applied to the deepest feature map $f_{4}$, Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
 $x_{\text{ASPP}}$ $=\phi_{\text{ASPP}}(f_{4})$
@@
 This module injects large-scale semantic context while retaining spatial discrimination.
-PSP [3]
-For the intermediate feature map $f_{3}$, Pyramid Scene Parsing aggregates multi-scale contextual cues by adaptive pooling at several grid scales:
+For the intermediate feature map $f_{3}$, Pyramid Scene Parsing i.e. PSP [21] aggregates multi-scale contextual cues by adaptive pooling at several grid scales:
 $x_{\text{PSP}}=\phi_{\text{PSP}}(f_{3})=\bigoplus_{s\in\{1,2,3,6\}}\sigma\!\left(\text{Up}\big(\text{Pool}_{s}(f_{3})\big)\right),$ (3)
@@
 III-D Cross-Attention Fusion
-The ASPP and PSP outputs are concatenated to form a joint query representation:
-$x_{\text{cat}}=[\,x_{\text{ASPP}}\,\|\,x_{\text{PSP}}\,]\in\mathbb{R}^{B\times(C_{a}+C_{p})\times H\times W}.$ (4)
-A convolutional multi-head cross-attention module then refines this fused tensor by treating $x_{\text{PSP}}$ as contextual guidance:
-$Q$ $=W_{Q}*x_{\text{cat}},$ (5)
-$K$ $=W_{K}*x_{\text{PSP}},$ (6)
-$V$ $=W_{V}*x_{\text{PSP}},$ (7)
+The outputs from the two multi-scale context extractors are concatenated to form a joint feature representation:
+$x_{\text{cat}}=[\,x_{\text{MC1}}\,\|\,x_{\text{MC2}}\,]\in\mathbb{R}^{B\times(C_{a}+C_{p})\times H\times W},$ (4)
+where $x_{\text{MC1}}$ emphasizes fine spatial detail from ASPP module and $x_{\text{MC2}}$ from PSP module encodes broader contextual semantics.
+To effectively combine these complementary cues, we employ a convolutional cross-attention mechanism.
+Here, $x_{\text{cat}}$ is used to generate the queries, as the concatenated representation benefits from selectively incorporating global context to refine local structure.
+Meanwhile, $x_{\text{MC2}}$ provides the keys and values, since its pooled multi-scale features offer stable, region-level guidance suitable for modulating finer details.
+This directional interaction allows each spatial location in $x_{\text{MC1}}$ to attend to the most relevant semantic context encoded by $x_{\text{MC2}}$.
+Formally, the attention is computed as follows: the query $Q=W_{Q}*x_{\text{cat}}$, key $K=W_{K}*x_{\text{MC2}}$, and value $V=W_{V}*x_{\text{MC2}}$ are first obtained.
 $\text{Attn}(Q,K,V)$ $=\operatorname*{softmax}\!\Big(\tfrac{QK^{\top}}{\sqrt{d_{k}}}\Big)V,$ (5)
-$\tilde{x}$ $=W_{O}*\text{Attn}(Q,K,V)+x_{\text{cat}}.$ (9)
-Here, * denotes convolution, $d_{k}$ is the per-head dimension, and residual addition ensures stable learning. This mechanism aligns global semantics from ASPP with fine-grained spatial cues from PSP.
+$\tilde{x}$ $=W_{O}*\text{Attn}(Q,K,V)+x_{\text{cat}},$ (6)
+where * denotes convolution and residual addition preserves the original spatial representation.
+This formulation enables fine-grained features to be adaptively enhanced by global contextual cues, producing a richer and more discriminative multi-scale representation.
 III-E Bottleneck Projection
@@
 $z^{\prime}=\phi_{\text{CA}}(z)=\phi_{\text{ECA}}(z)\odot\phi_{\text{SA}}(z)+z,$ (8)
-where $\phi_{\text{ECA}}$ denotes Efficient Channel Attention [23] and $\phi_{\text{SA}}$ is spatial attention [24, 25], and $\odot$ is element-wise multiplication.
-This highlights discriminative regions while preserving residual information.
+where $\phi_{\text{ECA}}$ denotes Efficient Channel Attention [22] and $\phi_{\text{SA}}$ denotes spatial attention [23, 24].
+This joint formulation first reweights channels based on their global importance and then emphasizes spatially salient regions, allowing the model to focus on both discriminative spectral responses and structurally meaningful areas.
 III-G Decoder with Deep Supervision
@@
 IV Experiments
-TABLE I: Performance comparison on CloudSEN12 L1C (TOA).
+TABLE I: Performance comparison on CloudSEN12 L2A (BOA).
 Model IoU (%) F1 (%) Accuracy (%) aAcc
 Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
-DBNet [15] 85.20 43.02 81.40 52.44 65.52 92.01 60.16 89.75 68.80 77.68 95.56 53.15 89.40 62.47 75.15 86.83
-CDNetv2 [13] 84.98 43.37 80.67 53.40 65.60 91.88 60.50 89.30 69.62 77.83 94.80 53.13 89.32 65.34 75.65 86.68
-HRCloudNet [18] 86.01 45.88 83.54 57.61 68.26 92.48 62.90 91.03 73.10 79.88 94.49 59.47 91.39 67.19 78.13 87.86
-U-Net-MobV2 90.82 85.98 50.54 67.21 73.64 95.19 92.46 67.14 80.39 83.80 94.42 96.11 95.05 96.62 95.55 91.10
-U-Net++-MobV2 90.93 86.58 49.78 66.93 73.56 95.25 92.81 66.47 80.19 83.68 94.47 96.25 95.27 96.56 95.64 91.28
-DeepLabV3+-MobV2 90.13 85.75 48.97 66.29 72.79 94.81 92.33 65.75 79.73 83.15 94.02 96.04 94.53 96.51 95.28 90.55
-U-Net-ResNet101 90.62 86.16 47.15 66.05 72.50 95.08 92.56 64.08 79.56 82.82 94.24 96.13 94.97 96.57 95.48 90.95
-U-Net++-ResNet101 90.90 86.43 46.24 68.46 73.01 95.23 92.72 63.24 81.27 83.12 94.45 96.16 95.06 96.74 95.60 91.20
-DeepLabV3+-ResNet101 90.65 85.66 50.60 66.36 73.32 95.10 92.28 67.20 79.78 83.59 94.38 95.94 94.94 96.47 95.43 90.87
-UPerNet-InternImage 90.85 85.28 36.70 68.87 70.42 95.09 91.82 47.94 80.98 78.96 94.50 96.43 95.24 96.85 95.75 91.51
-Mask2Former-DINOv2 88.50 82.10 43.61 63.62 69.46 93.90 90.17 60.73 77.76 80.64 92.94 94.84 94.41 96.11 94.57 89.15
-Mask2Former-Swin-T 90.75 86.70 49.90 69.53 74.22 95.15 92.88 66.58 82.02 84.16 94.46 96.32 94.34 96.93 95.51 91.03
-SegFormer-MiT-B5 91.47 86.70 49.77 70.24 74.54 95.55 92.88 66.46 82.52 84.35 94.84 96.27 95.25 96.92 95.82 91.64
-MSCloudCAM 91.64 87.24 52.58 70.63 75.52 95.64 93.19 68.92 82.78 85.13 94.93 96.42 95.58 97.04 95.99 91.99
+DBNet [12] 85.42 42.76 80.80 53.62 65.65 92.14 59.90 89.38 69.81 77.81 94.83 53.53 90.27 63.45 75.52 86.82
+CDNetv2 [10] 85.35 44.01 80.85 54.00 66.05 92.09 61.12 89.41 70.13 78.19 95.65 56.37 88.06 63.56 75.91 86.88
+HRCloudNet [15] 87.24 44.18 82.78 59.20 68.35 93.19 61.28 90.58 74.37 79.85 95.41 51.00 93.17 69.83 77.35 88.35
+U-Net-MobV2 88.74 83.87 54.13 60.15 71.72 94.03 91.23 70.24 75.12 82.65 93.45 95.55 93.71 95.43 94.53 89.07
+DeepLabV3+-MobV2 88.01 83.70 50.36 61.78 70.96 93.62 91.12 66.98 76.37 82.03 92.93 95.46 93.83 95.45 94.42 88.84
+DeepLabV3+-ResNet101 87.96 82.83 50.36 59.52 70.17 93.59 90.61 66.99 74.63 81.45 92.93 95.15 93.79 95.14 94.25 88.50
+UPerNet-InternImage 87.79 82.22 40.31 63.96 68.57 93.30 89.78 51.63 77.19 77.98 93.08 95.55 93.48 95.50 94.40 88.81
+Mask2Former-DINOv2 83.53 79.32 45.51 56.50 66.21 91.03 88.47 62.55 72.21 78.56 90.30 94.23 91.17 94.93 92.66 85.31
+Mask2Former-Swin-T 88.91 84.35 50.72 63.25 71.81 94.13 91.51 67.30 77.49 82.61 93.61 95.77 93.33 95.30 94.50 89.01
+SegFormer-MiT-B5 89.28 83.50 51.87 64.19 72.21 94.34 91.01 68.31 78.19 82.96 93.76 95.36 93.86 95.90 94.72 89.44
+MSCloudCAM 89.28 84.30 55.53 63.38 73.12 94.33 91.48 71.40 77.58 83.70 93.76 95.69 94.20 95.74 94.85 89.69
 IV-A Datasets
@@
 We use the high subset, selecting 10k samples each from L1C and L2A.
+L1C provides Top-of-Atmosphere (TOA) reflectance, while L2A provides Bottom-of-Atmosphere (BOA) surface reflectance.
 All bands are normalized by dividing top-of-atmosphere reflectance by $3000.0$.
@@
 IV-D Evaluation Metrics
-Per-class Metrics.
 $\mathrm{IoU}_{i}$ $=\frac{\mathrm{TP}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}},$ (12)
@@
 $\mathrm{Acc}_{i}$ $=\frac{\mathrm{TP}_{i}+\mathrm{TN}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}+\mathrm{TN}_{i}}.$ (14)
-Per-class metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes, giving equal weight regardless of class size.
+Class-wise metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes, giving equal weight regardless of class size.
 In contrast, overall accuracy (aAcc) is computed over all pixels, thus weighting larger classes more heavily.
-Ignored labels.
-Pixels with sentinel values (e.g., Fill in L8Biome) are removed from $\mathbf{C}$ and all denominators via an ignore_index mask, ensuring metrics are based on identical valid-pixel sets.
-Reporting.
-We report class-wise IoU and F1 (clear, thick cloud, thin cloud, cloud shadow), together with mIoU, mDice, mAcc, and aAcc. Unless noted, averages are macro over classes.
 V Results
 V-A Quantitative Performance
-We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables IIV present the detailed results.
-TABLE III: Performance comparison on CloudSEN12 L2A (BOA).
+We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables IIIIV present the detailed results.
+TABLE III: Performance comparison on CloudSEN12 L1C (TOA).
 Model IoU (%) F1 (%) Accuracy (%) aAcc
 Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
-DBNet [15] 85.42 42.76 80.80 53.62 65.65 92.14 59.90 89.38 69.81 77.81 94.83 53.53 90.27 63.45 75.52 86.82
-CDNetv2 [13] 85.35 44.01 80.85 54.00 66.05 92.09 61.12 89.41 70.13 78.19 95.65 56.37 88.06 63.56 75.91 86.88
-HRCloudNet [18] 87.24 44.18 82.78 59.20 68.35 93.19 61.28 90.58 74.37 79.85 95.41 51.00 93.17 69.83 77.35 88.35
-U-Net-MobV2 88.74 83.87 54.13 60.15 71.72 94.03 91.23 70.24 75.12 82.65 93.45 95.55 93.71 95.43 94.53 89.07
-U-Net++-MobV2 87.45 83.33 50.70 57.62 69.78 93.30 90.91 67.29 73.11 81.15 92.52 95.47 93.26 95.23 94.12 88.23
-DeepLabV3+-MobV2 88.01 83.70 50.36 61.78 70.96 93.62 91.12 66.98 76.37 82.03 92.93 95.46 93.83 95.45 94.42 88.84
-U-Net-ResNet101 86.51 81.76 37.12 58.89 66.07 92.77 89.97 54.14 74.12 77.75 91.83 94.73 92.76 95.13 93.61 87.23
-U-Net++-ResNet101 87.85 81.03 47.47 59.12 68.87 93.53 89.52 64.38 74.31 80.44 92.84 94.47 93.39 95.37 94.02 88.04
-DeepLabV3+-ResNet101 87.96 82.83 50.36 59.52 70.17 93.59 90.61 66.99 74.63 81.45 92.93 95.15 93.79 95.14 94.25 88.50
-UPerNet-InternImage 87.79 82.22 40.31 63.96 68.57 93.30 89.78 51.63 77.19 77.98 93.08 95.55 93.48 95.50 94.40 88.81
-Mask2Former-DINOv2 83.53 79.32 45.51 56.50 66.21 91.03 88.47 62.55 72.21 78.56 90.30 94.23 91.17 94.93 92.66 85.31
-Mask2Former-Swin-T 88.91 84.35 50.72 63.25 71.81 94.13 91.51 67.30 77.49 82.61 93.61 95.77 93.33 95.30 94.50 89.01
-SegFormer-MiT-B5 89.28 83.50 51.87 64.19 72.21 94.34 91.01 68.31 78.19 82.96 93.76 95.36 93.86 95.90 94.72 89.44
-MSCloudCAM 89.28 84.30 55.53 63.38 73.12 94.33 91.48 71.40 77.58 83.70 93.76 95.69 94.20 95.74 94.85 89.69
+DBNet [12] 85.20 43.02 81.40 52.44 65.52 92.01 60.16 89.75 68.80 77.68 95.56 53.15 89.40 62.47 75.15 86.83
+CDNetv2 [10] 84.98 43.37 80.67 53.40 65.60 91.88 60.50 89.30 69.62 77.83 94.80 53.13 89.32 65.34 75.65 86.68
+HRCloudNet [15] 86.01 45.88 83.54 57.61 68.26 92.48 62.90 91.03 73.10 79.88 94.49 59.47 91.39 67.19 78.13 87.86
+U-Net-MobV2 90.82 85.98 50.54 67.21 73.64 95.19 92.46 67.14 80.39 83.80 94.42 96.11 95.05 96.62 95.55 91.10
+DeepLabV3+-MobV2 90.13 85.75 48.97 66.29 72.79 94.81 92.33 65.75 79.73 83.15 94.02 96.04 94.53 96.51 95.28 90.55
+DeepLabV3+-ResNet101 90.65 85.66 50.60 66.36 73.32 95.10 92.28 67.20 79.78 83.59 94.38 95.94 94.94 96.47 95.43 90.87
+UPerNet-InternImage 90.85 85.28 36.70 68.87 70.42 95.09 91.82 47.94 80.98 78.96 94.50 96.43 95.24 96.85 95.75 91.51
+Mask2Former-DINOv2 88.50 82.10 43.61 63.62 69.46 93.90 90.17 60.73 77.76 80.64 92.94 94.84 94.41 96.11 94.57 89.15
+Mask2Former-Swin-T 90.75 86.70 49.90 69.53 74.22 95.15 92.88 66.58 82.02 84.16 94.46 96.32 94.34 96.93 95.51 91.03
+SegFormer-MiT-B5 91.47 86.70 49.77 70.24 74.54 95.55 92.88 66.46 82.52 84.35 94.84 96.27 95.25 96.92 95.82 91.64
+MSCloudCAM 91.64 87.24 52.58 70.63 75.52 95.64 93.19 68.92 82.78 85.13 94.93 96.42 95.58 97.04 95.99 91.99
 TABLE IV: Performance comparison on L8Biome.
@@
 U-Net-MobV2 88.98 78.59 39.97 6.37 53.48 94.17 88.01 57.11 11.98 62.82 94.26 91.50 87.55 98.55 92.97 85.93
-U-Net++-MobV2 88.96 81.68 39.44 4.42 53.62 94.16 89.91 56.57 8.47 62.28 94.13 93.55 88.76 98.57 93.55 87.09
 DeepLabV3+-MobV2 87.25 79.21 33.37 2.54 50.59 93.19 88.40 50.05 4.96 59.15 93.32 91.25 87.50 98.54 92.65 85.31
-U-Net-ResNet101 84.29 77.00 30.56 1.25 48.28 91.47 87.01 46.81 2.48 56.94 91.30 90.94 86.06 98.01 91.58 83.16
-U-Net++-ResNet101 88.16 80.23 34.94 3.81 51.78 93.71 89.03 51.78 7.34 60.46 93.73 91.94 87.87 98.52 93.01 86.03
 DeepLabV3+-ResNet101 87.56 74.19 34.24 4.88 50.22 93.37 85.18 51.01 9.31 59.72 93.47 89.56 85.51 98.56 91.78 83.55
@@
 V-B Qualitative Results
-To further illustrate the effectiveness of the proposed method, we present visual comparisons of cloud segmentation results in Fig.
-3.
-Each row corresponds to a different dataset, and each column corresponds to a different segmentation method.
-As shown in the figure, the proposed method produces more accurate delineation of thin clouds and cloud shadows, while reducing false detections compared to the baseline approaches.
+To further demonstrate the effectiveness of our method, Fig. 3 presents visual comparisons of cloud segmentation results. Each row shows a different dataset, and each column a different method.
 Figure 3: Qualitative cloud segmentation results.
@@
 Class–color mapping: Class 1 (Clear sky – light blue), Class 2 (Thick cloud – white), Class 3 (Thin cloud – light gray), Class 4 (Cloud shadow – dark gray).
-V-C Model Complexity Analysis
-Table V summarizes the computational complexity of all evaluated models in terms of FLOPs and parameter counts.
-While some architectures (e.g., UPerNet-InternImage and Mask2Former-DINOv2) exhibit significantly higher complexity, our proposed MSCloudCAM achieves a favorable trade-off between accuracy and efficiency.
-Specifically, MSCloudCAM attains competitive performance while requiring fewer FLOPs than SegFormer-MiT-B5, and substantially fewer FLOPs than UPerNet-InternImage, with a moderate parameter count.
-TABLE V: Model complexity comparison in terms of FLOPs and parameters.
-Model FLOPs (G) Parameters (M)
-U-Net-MobileNetV2 13.613 6.629
-U-Net++-MobileNetV2 18.023 6.825
-DeepLabV3+-MobileNetV2 6.057 4.379
-U-Net-ResNet101 62.258 51.514
-U-Net++-ResNet101 250.000 67.978
-DeepLabV3+-ResNet101 56.192 45.670
-UPerNet-InternImage-B-DCNv4 200–250 100–130
-Mask2Former-Swin-T 73.090 47.400
-Mask2Former-DINOv2-ViT-B/14 75–95 87–92
-SegFormer-MiT-B5 55.430 45.652
-MSCloudCAM 38.98 47.44
+V-C Ablation Study
+To evaluate the role of each multicontext component, we conduct ablation experiments by selectively replacing or removing modules.
+Replacing ASPP with PSP enhances global context but loses fine detail, while using ASPP instead of PSP improves multi-scale local features but sacrifices global scene understanding.
+Using PSP and ASPP together provides complementary global–local context and outperforms either module alone.
+Removing the Cross-Attention Module and using simple concatenation weakens the model’s ability to emphasize informative regions.
+The full MSCloudCAM system with PSP, ASPP, and CAM delivers the highest accuracy.
+TABLE V: Ablation study evaluating dual multi context extractor configurations and our cross-attention module (CAM) on CloudSEN12 L2A (BOA).
+Dual Extractor Fusion Type mIoU mDice mAcc aAcc
+PSP + PSP CAM 70.26 81.61 92.05 86.41
+ASPP + ASPP CAM 70.09 81.25 91.93 86.18
+PSP + ASPP Concat 71.48 82.44 93.52 88.02
+PSP + ASPP CAM 73.12 83.70 94.85 89.69
+V-D Model Complexity Analysis
+Table VI summarizes the computational complexity of all evaluated models in terms of FLOPs, parameter counts, and inference efficiency.
+In addition to latency (ms/image), we report throughput (TP, samples/s) to quantify batch inference performance.
+All experiments were conducted on a single NVIDIA A100 40GB GPU using PyTorch 2.3 and CUDA 12.1, with a batch size of 8 and an input resolution of 512×512.
+All models were evaluated in FP32 precision.
+Our proposed MSCloudCAM achieves a noteworthy balance between accuracy and computational efficiency and CAM module only adds 1.7M trainable parameters to the model.
+TABLE VI: Model complexity comparison in terms of FLOPs (G), parameters (M), inference time (ms/image), and TP as throughput (FPS).
+[…]
+U-Net-MobileNetV2 13.6 6.6 11.5 87.0
+DeepLabV3+-MobileNetV2 6.1 4.3 14.8 67.5
+DeepLabV3+-ResNet101 56.1 45.6 38.5 26.0
+UPerNet-InternImageDCNv4 193.5 110.8 52.0 19.2
+Mask2Former-Swin-T 73.09 47.4 33.6 29.7
+Mask2Former-DINOv2-B14 85.6 90.6 44.0 22.7
+SegFormer-MiT-B5 55.4 45.6 28.7 34.8
+MSCloudCAM 38.98 47.44 15.9 62.7
 VI Code and Model Availability
@@
 We presented MSCloudCAM, a cross attention with multi-scale context network for multispectral, multi-sensor cloud segmentation.
-By integrating transformer-based backbones with multi-scale context and attention mechanisms, our method achieves superior segmentation accuracy over strong baselines on CloudSEN12 and L8Biome datasets.
+By integrating hierarchical vision backbone with dual multi-scale context extractors and emlpoying our novel convolution based cross attention module, our method achieves superior segmentation accuracy over state-of-art-models on CloudSEN12 and L8Biome datasets.
 Future work will explore lightweight variants for onboard satellite processing and extend the model to spatiotemporal cloud tracking.

### v4

@@
 A key contribution of MSCloudCAM is the explicit modeling of multiple complementary multi-scale context extractors.
-And also, rather than simply stacking or concatenating their outputs, our formulation uses one extractor’s fine-resolution features and the other extractor’s global contextual representations enabling dynamic, scale-aware feature selection.
-Building on this idea, we design a new convolution-based cross attention adapter that effectively fuses localized, detailed information with broader multi-scale context.
+Our formulation uses one extractor’s fine-resolution features and the other extractor’s global contextual representations enabling dynamic, scale-aware feature selection.
+Building on this idea, we design a new convolution-based cross attention adaptation module (CAM) that effectively fuses localized, detailed information with broader multi-scale context rather than simply stacking or concatenating context extractors’ outputs.
 Integrated with a hierarchical vision backbone and refined through channel and spatial attention mechanisms, MSCloudCAM achieves strong spectral–spatial discrimination.
-Experiments on various multisensor datatsets e.g.
-CloudSEN12 (Sentinel-2) and L8Biome (Landsat-8) show that MSCloudCAM outperforms recent state-of-the-art models while maintaining competitive model complexity, highlighting the novelty and effectiveness of the proposed design for large-scale Earth observation.
+Experiments on multisensor datasets e.g.
+CloudSEN12 (Sentinel-2) and L8Biome (Landsat-8), demonstrate that MSCloudCAM achieves superior overall segmentation performance and competitive class-wise accuracy compared to recent state-of-the-art models, while maintaining efficient model complexity, highlighting the effectiveness of the proposed design for large-scale Earth observation.
+Index Terms—Cloud segmentation, Multispectral Imagery, Deep Learning, Convolutional Cross Attention, Multi-scale Context, Sentinel-2, Landsat-8.
 I Introduction
@@
 This limitation becomes especially critical in multi-class cloud segmentation requires modeling subtle spectral differences and large variations in spatial scale.
-Figure 1: High-level architecture of MSCloudCAM. The model integrates a hierarchical vision backbone with dual multi-scale context extractors and introduces a novel cross-attention module.
-To address these challenges, we propose MSCloudCAM, a cross-attention with multi-scale context network designed for robust cloud segmentation in multispectral, multi-sensor imagery.
-MSCloudCAM leverages complementary multi-scale context encoders and a newly designed cross-attention mechanism to dynamically fuse fine-grained spatial features with global contextual information.
-This design enables accurate pixel-wise classification across datasets with differing spectral responses, resolutions, and cloud characteristics.
+Figure 1: High-level architecture of MSCloudCAM.
+The model integrates a hierarchical vision backbone with dual multi-scale context extractors and introduces a novel convolutional cross-attention adaptation module (CAM).
+To address these challenges, we propose MSCloudCAM, a cross-attention adapter with multi-scale context network designed for robust cloud segmentation in multispectral, multi-sensor imagery.
 Our key contributions are as follows:
@@
 Thus it enables us to achieve comprehensive multi-scale feature utilization capability from a hierarchical vision backbone.
-  • We design a novel convolution-based cross-attention module, i.e.
+  • We design a novel convolution-based cross-attention adaptation module, i.e.
 CAM that leverages these complementary feature hierarchies through an adaptive formulation.
@@
 II-B Deep Learning Approaches
-Deep learning has become the dominant strategy for cloud segmentation due to its ability to model spatial–spectral structure directly from multispectral inputs.
-CNN-based architectures such as UNet [6] and UNetMobv2 [7] capture fine localization through skip connections.DeepLabV3+ [8] enhances multi-scale context via dilated convolutions and ASPP.
+CNN-based deep learning architectures such as UNet [6] and UNetMobv2 [7] capture fine localization through skip connections.DeepLabV3+ [8] enhances multi-scale context via dilated convolutions and ASPP.
 Numerous CNN variants target cloud detection specifically, including CDNetv1/v2 [9, 10], KappaMask [11], DBNet [12], SCNN [13], MCDNet [14], and HRCloudNet [15], each improving spectral–spatial fusion or multi-scale representation.
@@
 Despite significant progress, most approaches are designed for single-sensor datasets, limiting cross-sensor generalization.
-Moreover, few models explicitly integrate complementary multi-scale spectral–spatial cues using directional cross-attention.
+Moreover, few models explicitly integrate complementary multi-scale spectral–spatial cues using directional cross-attention adaptation.
 Our MSCloudCAM addresses both challenges.
@@
 MSCloudCAM is a multispectral semantic segmentation network that integrates a hierarchical vision backbone e.g.
-Swin Transformer [19] encoder with dual context extractor modules, a cross-attention fusion mechanism, and a lightweight combined attention refinement block.
+Swin Transformer [19] encoder with dual context extractor modules, a cross-attention adaptation mechanism, and a lightweight (only adds 1.7M trainable parameters) combined attention refinement block.
 The model is designed to capture long-range spectral–spatial dependencies while retaining fine-grained local structure.
@@
 ASPP and PSP modules enrich deep and intermediate representations.
-These are fused by convolutional cross-attention, refined via channel–spatial attention, and decoded by a multi-stage upsampling path with auxiliary supervision.
+These are fused by convolutional cross-attention adapter, refined via channel–spatial attention, and decoded by a multi-stage upsampling path with auxiliary supervision.
 III-B Backbone Feature Extraction
-Let $\mathbf{X}\in\mathbb{R}^{B\times C_{\text{in}}\times H_{0}\times W_{0}}$ denote the input multispectral image with $C_{\text{in}}$ spectral channels (13 for Sentinel-2, 11 for Landsat-8).
+Let $\mathbf{X}\in\mathbb{R}^{B\times C_{\text{in}}\times H_{0}\times W_{0}}$ denote the input multispectral image with $C_{\text{in}}$ spectral channels (13 for Sentinel-2, 11 for Landsat-8), where $B$ is the batch size and $(H_{0},W_{0})$ denote the input spatial resolution.
 The encoder $\mathcal{E}(\cdot)$ produces a hierarchy of features:
 $\{f_{1},f_{2},f_{3},f_{4}\}=\mathcal{E}(\mathbf{X}),\quad f_{i}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}},$ (1)
-with channel dimensions $[96,192,384,768]$ and progressively reduced spatial sizes $(H_{i},W_{i})$.
-Shifted-window self-attention within Swin ensures that $f_{4}$ captures global dependencies, while $f_{1}$$f_{3}$ preserve finer detail.
+where $(H_{i},W_{i})$ denote the spatial resolutions at stage $i$, progressively reduced from $(H_{0},W_{0})$ through hierarchical downsampling, and $C_{i}\in\{96,192,384,768\}$.
+Shifted-window self-attention within Swin ensures that $f_{4}$ captures global dependencies, while $f_{1}$$f_{3}$ preserve finer spatial details.
 III-C Multi-Scale Context Modules
 To enhance representational richness, we employ two complementary multi context encoders that operate at different feature hierarchies.
-As a part of dual multiscale context extractor, ASPP [8] is applied to the deepest feature map $f_{4}$, Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
+As part of the dual multiscale context extractor, ASPP [8] is applied to the deepest feature map $f_{4}$.
+Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
 $x_{\text{ASPP}}$ $=\phi_{\text{ASPP}}(f_{4})$
@@
 where $\oplus$ denotes channel-wise concatenation, $r$ specifies the dilation rate, GAP is global average pooling, and $\sigma$ represents the ReLU nonlinearity.
-This module injects large-scale semantic context while retaining spatial discrimination.
-For the intermediate feature map $f_{3}$, Pyramid Scene Parsing i.e. PSP [21] aggregates multi-scale contextual cues by adaptive pooling at several grid scales:
+This operation aggregates contextual information across multiple receptive fields.
+For the intermediate feature map $f_{3}$, Pyramid Scene Parsing (PSP) [21] aggregates contextual cues by adaptive pooling at several grid scales:
 $x_{\text{PSP}}=\phi_{\text{PSP}}(f_{3})=\bigoplus_{s\in\{1,2,3,6\}}\sigma\!\left(\text{Up}\big(\text{Pool}_{s}(f_{3})\big)\right),$ (3)
-where $\text{Pool}_{s}$ denotes average pooling to an $s\times s$ grid, followed by projection and bilinear upsampling.
-This captures scene-level information at progressively coarser scales, which is essential for delineating fine structures such as thin clouds.
-III-D Cross-Attention Fusion
+where where $s\in\{1,2,3,6\}$ denotes the pooling grid size and $\text{Pool}_{s}$ maps $f_{3}$ to an $s\times s$ grid via average pooling, followed by projection and bilinear upsampling.
+This operation encodes multi-scale scene layout information.
+This asymmetric placement aligns contextual operators with the feature hierarchy.
+PSP is applied to $f_{3}$ to exploit its higher spatial resolution for preserving structural details and accurate boundary localization, whereas ASPP is applied to the deeper $f_{4}$ where stronger semantic abstractions and a coarser resolution benefit more from large-dilation convolutions and global context aggregation.
+III-D Cross-Attention Adaptation
 The outputs from the two multi-scale context extractors are concatenated to form a joint feature representation:
@@
 where $x_{\text{MC1}}$ emphasizes fine spatial detail from ASPP module and $x_{\text{MC2}}$ from PSP module encodes broader contextual semantics.
-To effectively combine these complementary cues, we employ a convolutional cross-attention mechanism.
+To effectively combine these complementary cues, we employ a convolutional cross-attention adaptation mechanism.
 Here, $x_{\text{cat}}$ is used to generate the queries, as the concatenated representation benefits from selectively incorporating global context to refine local structure.
@@
 where $Y$ is the ground truth mask, and $(\lambda_{\text{final}},\lambda_{1},\lambda_{2})$ are supervision weights.
-IV Experiments
-TABLE I: Performance comparison on CloudSEN12 L2A (BOA).
+TABLE I: Performance comparison on CloudSEN12 L1C (TOA).
+Model IoU (%) F1 (%) Accuracy (%) aAcc
+Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
+DBNet [12] 85.20 43.02 81.40 52.44 65.52 92.01 60.16 89.75 68.80 77.68 95.56 53.15 89.40 62.47 75.15 86.83
+CDNetv2 [10] 84.98 43.37 80.67 53.40 65.60 91.88 60.50 89.30 69.62 77.83 94.80 53.13 89.32 65.34 75.65 86.68
+HRCloudNet [15] 86.01 45.88 83.54 57.61 68.26 92.48 62.90 91.03 73.10 79.88 94.49 59.47 91.39 67.19 78.13 87.86
+UNet-MobV2 90.82 85.98 50.54 67.21 73.64 95.19 92.46 67.14 80.39 83.80 94.42 96.11 95.05 96.62 95.55 91.10
+DeepLabV3+-MobV2 90.13 85.75 48.97 66.29 72.79 94.81 92.33 65.75 79.73 83.15 94.02 96.04 94.53 96.51 95.28 90.55
+UPerNet-InternImage 90.85 85.28 36.70 68.87 70.42 95.09 91.82 47.94 80.98 78.96 94.50 96.43 95.24 96.85 95.75 91.51
+Mask2Former-DINOv2 88.50 82.10 43.61 63.62 69.46 93.90 90.17 60.73 77.76 80.64 92.94 94.84 94.41 96.11 94.57 89.15
+Mask2Former-Swin-T 90.75 86.70 49.90 69.53 74.22 95.15 92.88 66.58 82.02 84.16 94.46 96.32 94.34 96.93 95.51 91.03
+SegFormer-MiT-B5 91.47 86.70 49.77 70.24 74.54 95.55 92.88 66.46 82.52 84.35 94.84 96.27 95.25 96.92 95.82 91.64
+MSCloudCAM 91.64 87.24 52.58 70.63 75.52 95.64 93.19 68.92 82.78 85.13 94.93 96.42 95.58 97.04 95.99 91.99
+TABLE II: Performance comparison on CloudSEN12 L2A (BOA).
 Model IoU (%) F1 (%) Accuracy (%) aAcc
@@
 HRCloudNet [15] 87.24 44.18 82.78 59.20 68.35 93.19 61.28 90.58 74.37 79.85 95.41 51.00 93.17 69.83 77.35 88.35
-U-Net-MobV2 88.74 83.87 54.13 60.15 71.72 94.03 91.23 70.24 75.12 82.65 93.45 95.55 93.71 95.43 94.53 89.07
+UNet-MobV2 88.74 83.87 54.13 60.15 71.72 94.03 91.23 70.24 75.12 82.65 93.45 95.55 93.71 95.43 94.53 89.07
 DeepLabV3+-MobV2 88.01 83.70 50.36 61.78 70.96 93.62 91.12 66.98 76.37 82.03 92.93 95.46 93.83 95.45 94.42 88.84
-DeepLabV3+-ResNet101 87.96 82.83 50.36 59.52 70.17 93.59 90.61 66.99 74.63 81.45 92.93 95.15 93.79 95.14 94.25 88.50
 UPerNet-InternImage 87.79 82.22 40.31 63.96 68.57 93.30 89.78 51.63 77.19 77.98 93.08 95.55 93.48 95.50 94.40 88.81
@@
 MSCloudCAM 89.28 84.30 55.53 63.38 73.12 94.33 91.48 71.40 77.58 83.70 93.76 95.69 94.20 95.74 94.85 89.69
-IV-A Datasets
+III-H Datasets
 We evaluate MSCloud on two publicly available multispectral cloud segmentation datasets.
 CloudSEN12 [1] is a global Sentinel-2 cloud segmentation dataset with 13 spectral bands (10–60 m) and pixel-level labels for clear (0), thick cloud (1), thin cloud (2), and cloud shadow (3).
-We use the high subset, selecting 10k samples each from L1C and L2A.
-L1C provides Top-of-Atmosphere (TOA) reflectance, while L2A provides Bottom-of-Atmosphere (BOA) surface reflectance.
+We use the high subset, selecting 10k samples each from L1C (Top-of-Atmosphere - TOA) and L2A ( Bottom-of-Atmosphere - BOA).
 All bands are normalized by dividing top-of-atmosphere reflectance by $3000.0$.
-L8Biome [2] is derived from Landsat-8 CCA, containing 11 bands (30 m) with manual masks.
+L8Biome [2] is derived from Landsat-8 Cloud Cover Assessment, containing 11 bands (30 m) with manual masks.
 The original five classes—fill (0), shadow (64), clear (128), thin cloud (192), and cloud (255)—are remapped to CloudSEN12’s taxonomy: clear $\rightarrow 0$, thick cloud $\rightarrow 1$, thin cloud $\rightarrow 2$, shadow $\rightarrow 3$.
@@
 Bands are normalized with the same $3000.0$ divisor for consistency.
-TABLE II: Dataset split statistics for CloudSEN12 and L8Biome.
+TABLE III: Dataset split statistics for CloudSEN12 and L8Biome.
 Dataset Train Validation Test
@@
 L8Biome 13,378 4,459 4,459
-IV-B Loss Function
+III-I Loss Function
 The training objective is a weighted sum of the losses from the final prediction and the two auxiliary predictions:
@@
 We set $\lambda_{\text{final}}=1.0$ and $\lambda_{1}=\lambda_{2}=0.4$, ensuring the final prediction dominates while auxiliary supervision stabilizes early-stage training.
-IV-C Training Protocol
+III-J Training Protocol
 Our model is trained using the Adam optimizer with an initial learning rate of $1\times 10^{-4}$. The batch size is set to 8, and we train for 100 epochs.
-IV-D Evaluation Metrics
+III-K Evaluation Metrics
 $\mathrm{IoU}_{i}$ $=\frac{\mathrm{TP}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}},$ (12)
@@
 $\mathrm{Acc}_{i}$ $=\frac{\mathrm{TP}_{i}+\mathrm{TN}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}+\mathrm{TN}_{i}}.$ (14)
-Class-wise metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes, giving equal weight regardless of class size.
-In contrast, overall accuracy (aAcc) is computed over all pixels, thus weighting larger classes more heavily.
-V Results
-V-A Quantitative Performance
-We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables IIIIV present the detailed results.
-TABLE III: Performance comparison on CloudSEN12 L1C (TOA).
-Model IoU (%) F1 (%) Accuracy (%) aAcc
-Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
-DBNet [12] 85.20 43.02 81.40 52.44 65.52 92.01 60.16 89.75 68.80 77.68 95.56 53.15 89.40 62.47 75.15 86.83
-CDNetv2 [10] 84.98 43.37 80.67 53.40 65.60 91.88 60.50 89.30 69.62 77.83 94.80 53.13 89.32 65.34 75.65 86.68
-HRCloudNet [15] 86.01 45.88 83.54 57.61 68.26 92.48 62.90 91.03 73.10 79.88 94.49 59.47 91.39 67.19 78.13 87.86
-U-Net-MobV2 90.82 85.98 50.54 67.21 73.64 95.19 92.46 67.14 80.39 83.80 94.42 96.11 95.05 96.62 95.55 91.10
-DeepLabV3+-MobV2 90.13 85.75 48.97 66.29 72.79 94.81 92.33 65.75 79.73 83.15 94.02 96.04 94.53 96.51 95.28 90.55
-DeepLabV3+-ResNet101 90.65 85.66 50.60 66.36 73.32 95.10 92.28 67.20 79.78 83.59 94.38 95.94 94.94 96.47 95.43 90.87
-UPerNet-InternImage 90.85 85.28 36.70 68.87 70.42 95.09 91.82 47.94 80.98 78.96 94.50 96.43 95.24 96.85 95.75 91.51
-Mask2Former-DINOv2 88.50 82.10 43.61 63.62 69.46 93.90 90.17 60.73 77.76 80.64 92.94 94.84 94.41 96.11 94.57 89.15
-Mask2Former-Swin-T 90.75 86.70 49.90 69.53 74.22 95.15 92.88 66.58 82.02 84.16 94.46 96.32 94.34 96.93 95.51 91.03
-SegFormer-MiT-B5 91.47 86.70 49.77 70.24 74.54 95.55 92.88 66.46 82.52 84.35 94.84 96.27 95.25 96.92 95.82 91.64
-MSCloudCAM 91.64 87.24 52.58 70.63 75.52 95.64 93.19 68.92 82.78 85.13 94.93 96.42 95.58 97.04 95.99 91.99
+Here, $\mathrm{TP}_{i}$, $\mathrm{FP}_{i}$, $\mathrm{FN}_{i}$, and $\mathrm{TN}_{i}$ denote the numbers of true positives, false positives, false negatives, and true negatives for class $i$, respectively.
+Class-wise metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes.
+In contrast, overall accuracy (aAcc) is computed over all pixels.
+IV Results
+IV-A Quantitative Performance
+We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables I, II, IV present the detailed results.
 TABLE IV: Performance comparison on L8Biome.
@@
 Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
+DBNet [12] 80.19 38.18 82.37 04.90 51.41 89.00 55.25 90.34 09.34 60.99 96.51 49.34 87.89 05.03 59.70 83.62
 CDNetv2 [10] 73.81 24.85 75.85 00.00 43.63 84.93 39.81 86.27 00.00 52.75 90.18 31.95 89.81 00.00 52.98 78.56
 HRCloudNet [15] 72.41 30.89 70.76 00.00 43.51 83.99 47.20 82.88 00.00 53.52 85.50 43.70 85.88 00.00 53.77 77.04
-DBNet [12] 80.19 38.18 82.37 04.90 51.41 89.00 55.25 90.34 09.34 60.99 96.51 49.34 87.89 05.03 59.70 83.62
-U-Net-MobV2 88.98 78.59 39.97 6.37 53.48 94.17 88.01 57.11 11.98 62.82 94.26 91.50 87.55 98.55 92.97 85.93
+UNet-MobV2 88.98 78.59 39.97 6.37 53.48 94.17 88.01 57.11 11.98 62.82 94.26 91.50 87.55 98.55 92.97 85.93
 DeepLabV3+-MobV2 87.25 79.21 33.37 2.54 50.59 93.19 88.40 50.05 4.96 59.15 93.32 91.25 87.50 98.54 92.65 85.31
-DeepLabV3+-ResNet101 87.56 74.19 34.24 4.88 50.22 93.37 85.18 51.01 9.31 59.72 93.47 89.56 85.51 98.56 91.78 83.55
 UPerNet-InternImage 54.68 46.97 18.65 1.56 30.46 60.19 53.40 26.72 2.55 35.72 91.00 90.58 83.73 98.35 90.92 81.83
@@
 MSCloudCAM 88.27 82.23 40.19 21.48 58.04 93.77 90.25 57.34 35.36 69.18 93.84 92.91 88.74 98.70 93.55 87.09
-V-B Qualitative Results
-To further demonstrate the effectiveness of our method, Fig. 3 presents visual comparisons of cloud segmentation results. Each row shows a different dataset, and each column a different method.
-Figure 3: Qualitative cloud segmentation results.
-Each row corresponds to a dataset.
-The proposed method produces sharper delineation of thin clouds and shadows compared to other approaches.
-Class–color mapping: Class 1 (Clear sky – light blue), Class 2 (Thick cloud – white), Class 3 (Thin cloud – light gray), Class 4 (Cloud shadow – dark gray).
-V-C Ablation Study
+IV-B Qualitative Results
+To further demonstrate the effectiveness of our method, Fig. 3 presents visual comparisons of cloud segmentation results.
+Figure 3: Qualitative cloud segmentation results across datasets.
+Each row represents a dataset.The proposed method shows sharper delineation of thin clouds and shadows than other approaches.
+Class–color mapping: Class 1 (Clear sky – light blue), Class 2 (Thick cloud – white), Class 3 (Thin cloud – light gray), Class 4 (Cloud shadow – dark gray)
+IV-C Ablation Study
 To evaluate the role of each multicontext component, we conduct ablation experiments by selectively replacing or removing modules.
@@
 Using PSP and ASPP together provides complementary global–local context and outperforms either module alone.
-Removing the Cross-Attention Module and using simple concatenation weakens the model’s ability to emphasize informative regions.
+Removing the cross-attention module (CAM) and using simple concatenation weakens the model’s ability to emphasize informative regions.
 The full MSCloudCAM system with PSP, ASPP, and CAM delivers the highest accuracy.
-TABLE V: Ablation study evaluating dual multi context extractor configurations and our cross-attention module (CAM) on CloudSEN12 L2A (BOA).
-Dual Extractor Fusion Type mIoU mDice mAcc aAcc
-PSP + PSP CAM 70.26 81.61 92.05 86.41
-ASPP + ASPP CAM 70.09 81.25 91.93 86.18
-PSP + ASPP Concat 71.48 82.44 93.52 88.02
-PSP + ASPP CAM 73.12 83.70 94.85 89.69
-V-D Model Complexity Analysis
-Table VI summarizes the computational complexity of all evaluated models in terms of FLOPs, parameter counts, and inference efficiency.
+Overall, these findings confirm that each component contributes uniquely to the system.
+TABLE V: Ablation study evaluating dual multi context extractor and cross-attention module (CAM) on CloudSEN12 L1C, CloudSEN L2A, L8Biome datasets.
+Dataset Extractor Fusion mIoU mF1 mAcc
+CloudSEN12 PSP + PSP CAM 72.31 83.56 94.11
+L1C ASPP + ASPP CAM 71.98 83.37 93.89
+PSP + ASPP Concat 73.43 84.29 94.75
+PSP + ASPP CAM 75.52 85.13 95.99
+CloudSEN12 PSP + PSP CAM 70.26 81.61 92.05
+L2A ASPP + ASPP CAM 70.09 81.25 91.92
+PSP + ASPP Concat 71.48 82.44 92.63
+PSP + ASPP CAM 73.12 83.70 94.85
+L8Biome PSP + PSP CAM 54.95 67.63 91.87
+ASPP + ASPP CAM 55.08 67.31 91.46
+PSP + ASPP Concat 56.41 68.22 92.02
+PSP + ASPP CAM 58.04 69.18 93.55
+IV-D Model Complexity Analysis
+Table VI summarizes the computational complexity of evaluated models in terms of FLOPs, parameter counts, and inference efficiency.
 In addition to latency (ms/image), we report throughput (TP, samples/s) to quantify batch inference performance.
@@
 All models were evaluated in FP32 precision.
-Our proposed MSCloudCAM achieves a noteworthy balance between accuracy and computational efficiency and CAM module only adds 1.7M trainable parameters to the model.
+Our proposed MSCloudCAM achieves a noteworthy balance between accuracy and computational efficiency and CAM only adds 1.7M trainable parameters to the model.
 TABLE VI: Model complexity comparison in terms of FLOPs (G), parameters (M), inference time (ms/image), and TP as throughput (FPS).
@@
 U-Net-MobileNetV2 13.6 6.6 11.5 87.0
-DeepLabV3+-MobileNetV2 6.1 4.3 14.8 67.5
+DeepLabV3+-MobV2 6.1 4.3 14.8 67.5
 DeepLabV3+-ResNet101 56.1 45.6 38.5 26.0
-UPerNet-InternImageDCNv4 193.5 110.8 52.0 19.2
+UPerNet-InternImage 193 110 52.0 19.2
 Mask2Former-Swin-T 73.09 47.4 33.6 29.7
-Mask2Former-DINOv2-B14 85.6 90.6 44.0 22.7
+Mask2Former-DINOv2 85.6 90.6 44.0 22.7
 SegFormer-MiT-B5 55.4 45.6 28.7 34.8
 MSCloudCAM 38.98 47.44 15.9 62.7
-VI Code and Model Availability
-The full implementation of our proposed MS-CloudCAM framework is publicly available.
+V Code and Model Availability
+The full implementation of our proposed MSCloudCAM framework is publicly available.
 The source code, training scripts, and experimental configurations are hosted on GitHub: https://github.com/mazid-rafee/ms-cloudcam.
 Additionally, pretrained model checkpoints are released on Hugging Face: https://huggingface.co/mazid-rafee/MS-CloudCAM/tree/main.
-VII Conclusion
-We presented MSCloudCAM, a cross attention with multi-scale context network for multispectral, multi-sensor cloud segmentation.
-By integrating hierarchical vision backbone with dual multi-scale context extractors and emlpoying our novel convolution based cross attention module, our method achieves superior segmentation accuracy over state-of-art-models on CloudSEN12 and L8Biome datasets.
-Future work will explore lightweight variants for onboard satellite processing and extend the model to spatiotemporal cloud tracking.
+VI Conclusion
+We presented MSCloudCAM which integrates hierarchical vision backbone with dual multi-scale context extractors and employs our novel convolution based cross attention adaptation module.
+Extensive experiments on the CloudSEN12 and L8Biome datasets demonstrate that MSCloudCAM consistently surpasses state-of-the-art methods across multiple segmentation metrics.
+These highlight the effectiveness and generalizability of our design for multi-sensor and multi-spectral cloud understanding.
+While our experiments focus on transformer-based backbones, evaluating the proposed context adapter with CNN-based encoders remains an interesting direction for future work.
 Acknowledgment